# thanks: These two authors contributed equally

### v1

@@ III.2 Single emitter, illumination through the waveguide
Now let us consider the case where the emitter is driven through the waveguide with Rabi frequency $\Omega_{\mathrm{in}}^{\mathrm{wg}}$, see Fig. 1, and a corresponding amplitude of the input coherent state $\alpha_{\mathrm{in}}=\Omega^{\mathrm{wg}}_{\mathrm{in}}/(2\sqrt{\beta\Gamma})$. The coherent component at the waveguide output $\alpha_{\mathrm{out}}$ now consists of the coherently scattered light $\alpha_{\mathrm{sc}}$ that interferes with the coherent field that is already present inside the waveguide $\alpha_{\mathrm{in}}$:
@@
In the limit of $\beta\ll 1$, the scattered field of a single atom is much smaller than the input field, $\alpha_{\mathrm{sc}}\ll\alpha_{\mathrm{in}}$, such that the two-photon wavefunction at the output is identical to the one of the input, $\psi(\tau)\approx\alpha_{\mathrm{in}}^{2}$. Thus, a single weakly coupled emitter has a negligible effect on the photon statistics of the light transmitted in the waveguide. In contrast, as we discuss in the following, the interaction of $N$ weakly coupled emitters, can strongly suppress or enhance the two-photon wavefunction at zero time delay $\psi(\tau=0)$.
@@ III.3 Single emitter, combined illumination
Let us now consider the case where the emitter is driven by both an external field and a guided field with Rabi frequencies $\Omega^{\mathrm{ext}}$ and $\Omega_{\mathrm{in}}^{\mathrm{wg}}$, respectively. Here and in the following, when considering combined illumination, for simplicity we assume identical polarizations and detunings for both fields, such that the Rabi frequencies sum up to the total Rabi frequency seen by the atom $\Omega=\Omega^{\mathrm{ext}}+\Omega_{\mathrm{in}}^{\mathrm{wg}}$. The combined scattering rate is then
@@
where we introduced the effective coupling constant $\beta^{\prime}=\beta\left(1+\Omega^{\mathrm{ext}}/\Omega_{\mathrm{in}}^{\mathrm{wg}}\right)$ which depends on the ratio of the driving intensity $\Omega_{\mathrm{in}}^{\mathrm{wg}}$ and the intensity resulting from the external Bragg beam $\Omega^{\mathrm{ext}}$. In analogy to Sec. III.2, one can introduce an effective transmission $t^{\prime}_{\Delta}$ and normalized scattering coefficient $g^{\prime}_{\Delta}$
@@
In summary, under a combined illumination, a single emitter effectively behaves, as if illuminated through the waveguide with an effectively increased coupling constant $\beta^{\prime}$, as discussed in details in [67].
@@ IV Many emitters
Here, the coherent scattering amplitudes $\alpha_{n}^{\textrm{sc}}$ are computed according to Eq. (7) using the local Rabi frequencies $\Omega_{n}$ seen by the atoms, which depend on the illumination configuration, see below. As previously, the two-photon coherent amplitude then is given by $\psi_{\mathrm{coh}}=\alpha_{\mathrm{out}}^{2}$.
For deriving the incoherent two-photon amplitude at the waveguide output, we resort to the frequency domain and find:
@@
where $\psi_{\mathrm{incoh}}^{(n)}(\omega)$ is the frequency-domain presentation of the incoherent component of the two-photon amplitude. We calculate it according to Eq. (10) using the local Rabi frequency $\Omega_{n}$, thereby assuming that the incoherent scattering process is only driven by the coherent component of the incident field. The product of amplitude transmission coefficients, $t_{\Delta+\omega}t_{\Delta-\omega}$, gives the amplitude transmission coefficient of a pair of incoherently scattered blue- and red-detuned photons past one atom. Note that, here, we neglect higher order nonlinear interaction where incoherently scattered photon pairs experience a second non-linear scattering. Neglecting these higher-order incoherent scattering processes is a good approximation in the regime of weak coupling ($\beta\ll 1$), see Sec. IV.1 where we explicitly compare our result with [33].
[…]
Figure 4: a) Contribution of the $n$-th atom to the incoherent two-photon component of the transmitted light. The first $n-1$ emitters attenuate the input light which is then incoherently scattered by the $n$-th emitter. Then, the following $N-n$ emitters act as a dispersive and absorptive medium for this incoherent two-photon component. b) The evolution of the coherent amplitude $\psi_{\mathrm{coh}}$ and incoherent photon amplitude $\psi_{\mathrm{incoh}}(\tau=0)$ at zero time delay as a function of atom number for two different detunings $\Delta$. The amplitudes are normalized to $\alpha_{\mathrm{in}}$. c) Second-order correlation function at zero time delay, $g^{(2)}(0)$, for different detunings. Plots are shown for $\beta=0.007$ as in [3].
@@ IV.1 N emitters, illumination through the waveguide
First, we discuss the case where $N$ atoms are solely excited through the waveguide mode. This setting was theoretically studied in [33] and [60] based on a scattering matrix formalism and higher order cumulant expansion, respectively, with a detailed discussion of the resonant excitation case ($\Delta=0$). Here, we extend this discussion to off-resonant excitation, which has been experimentally studied and theoretically analyzed [3] using the model presented here.
@@
Inserting this expression into the incoherent two-photon component emitted by the $n$-th emitter into the waveguide, $\psi^{(n)}_{\mathrm{incoh}}(\omega)$:
@@
Summing the incoherent contribution of all emitters, we obtain the incoherent component of the output two-photon wavefunction:
$\psi_{\mathrm{incoh}}(\omega)=(\Omega_{\mathrm{in}}^{\mathrm{wd}})^{2}\sum_{n=1}^{N}t_{\Delta}^{2(n-1)}\tilde{\psi}_{\mathrm{incoh}}(\omega)\left(t_{\Delta+\omega}t_{\Delta-\omega}\right)^{N-n}$
$=(\Omega_{\mathrm{in}}^{\mathrm{wd}})^{2}\,\tilde{\psi}_{\mathrm{incoh}}(\omega)\times\begin{cases}\frac{t_{\Delta}^{2N}-\left[t_{\Delta+\omega}t_{\Delta-\omega}\right]^{N}}{t_{\Delta}^{2}-t_{\Delta+\omega}t_{\Delta-\omega}}&\omega\neq 0\\ Nt_{\Delta}^{2(N-1)}&\omega=0\end{cases}$ (26)
@@
$\tilde{\psi}_{\mathrm{incoh}}(\omega)=\psi^{(n)}_{\mathrm{incoh}}/\Omega_{n}^{2}~{}.$ (27)
Thus, the incoherent component of the output two-photon wavefunction $\psi_{\mathrm{incoh}}(\omega)$ results from an interplay of (i) attenuation of coherent light along the ensemble, (ii) incoherent scattering at each atom and (iii) absorption of the incoherent photons by the following atoms, as sketched in Fig. 4.a). From the total two-photon wavefunction at the output, $\psi(\tau)$ , we obtain the second-order correlation function $g^{(2)}(\tau)$:
@@
Here, as $\psi_{\mathrm{incoh}}\propto\alpha_{\mathrm{in}}^{2}$, $g^{(2)}(\tau)$ is independent of the input power. This is a consequence of the truncation to the two-photon number state used here  [^5] Our theory is in leading order in $|\Omega|/\Gamma$ as in [33]. Correction to our model are on the order $P_{\mathrm{in}}\propto|\Omega|^{2}$ for $g^{2}(\tau)$, and $P_{\mathrm{}}^{2}$ for the squeezing spectrum.
@@
In this context, it is useful to define the phase difference between the total incoherent and coherent component at zero time delay, $\Delta\varphi=\mathrm{Arg}\{\psi_{\mathrm{incoh}}(\tau=0)\}-\mathrm{Arg}\{\psi_{\mathrm{coh}}\}$. Depending on $\Delta\varphi$, the incoherent two-photon component will interfere with the coherent component either destructively (when $\Delta\varphi=\left(2m+1\right)\pi,\,m\in\mathbb{Z}$) or constructively (when $\Delta\varphi=2m\pi,\,m\in\mathbb{Z}$). This is illustrated in Fig. 4 b) where we show the evolution of both the coherent and incoherent two-photon amplitude in the complex plane as a function of the number of emitters for two different detunings. The coherent component still decays exponentially with the number of emitters, as in the case for resonant excitation. By comparison, the incoherent two-photon amplitude at zero time delay $\psi_{\mathrm{incoh}}(\tau=0)$ shows a more complex dynamics. Its magnitude first linearly increases due to collective enhancement in forward scattering  [^6] See Appendix B and [2] for more details.. For larger atom number, the magnitude of the two-photon component decreases sub-exponentially with $N$ [33].
@@
Turning now to Fig. 4.c), it is apparent that the combined dynamics of $\psi_{\mathrm{coh}}$ and $\psi_{\mathrm{incoh}}(\tau=0)$ leads to an oscillatory behavior of $g^{(2)}(0)$ as a function of $N$, which stems from the $N$-dependence of $\Delta\varphi$, see panel b), leading to alternating constructive and destructive interference of the two components. Surprisingly, when scanning $N$ and $\Delta$, we find that several combinations lead to perfect destructive interference and, hence, to photon antibunching (see orange and green lines in panel c). Specifically, this occurs when the moduli of both components are equal, $|\psi_{\mathrm{coh}}|=|\psi_{\mathrm{incoh}}(\tau=0)|$, and their phase difference is $\Delta\varphi=(2m+1)\pi$, $m\in\mathbb{Z}$. While the antibunching point on resonance ($\Delta=0$) has been predicted and observed in [33, 1], perfect antibunching from off-resonant excitations is a new prediction. It has been recently observed experimentally based on the use of this model [3].
We now compare our results to Mahmoodian et al. [33] where the two-photon wavefunction has been calculated using a scattering formalism together with the Bethe-Ansatz technique [63]. In contrast to our model, this approach holds for any value of $\beta$. In Fig. 5 we compare the incoherent component $|\psi_{\mathrm{incoh}}(\tau=0)|$ calculated with the two approaches. Both predictions agree very well for $\beta\ll 1$. For larger $\beta$ deviations are expected as our model does not take into account higher order scattering processes. Note that, in contrast to [33], our model provides a closed expression for $\psi_{\mathrm{incoh}}(\tau=0)$, which is given by the frequency integral of the analytical expression given in Eq. (26). Thus, its numerical complexity does not depend on $N$, contrary to the numerical calculation based on [33]. Consequently, predictions made with our model are numerically much less expensive for large $N$.
@@ IV.2 External illumination: Bragg angle
where $\lambda_{f}=\lambda/n_{\mathrm{eff}}$ is the wavelength inside the waveguide, $n_{\mathrm{eff}}$ is the effective refractive index of the guided mode, and $a$ the lattice spacing and $m$ the Bragg order. This situation resembles the situation studied in [12, 72]. We will neglect free space interaction between the atoms such that we only consider Bragg conditions fulfilling $a\gtrsim\lambda$. We note that possible voids in the lattice would not affect our predictions.
For a plane wave, each emitter is subject to an identical external driving field with a Rabi frequency of magniture $|\Omega^{\mathrm{ext}}|$. The first emitter scatters a coherent amplitude $\alpha_{1}^{\mathrm{sc}}$ into the waveguide which constructively interferes with the scattered amplitude from the second emitter according to the Bragg condition of Eq. (29). After $N$ emitters, and taking into account absorption, the output two-photon coherent amplitude inside the waveguide is
$\alpha_{\mathrm{out}}$ $=\sum_{n=1}^{N}t_{0}^{N-n}\alpha_{1}^{\mathrm{sc}}=\frac{1-t_{0}^{N}}{1-t_{0}}\alpha_{1}^{\mathrm{sc}}=\frac{1-t_{0}^{N}}{2\beta}\alpha_{1}^{\mathrm{sc}}$ (30)
@@
using $t_{-\omega}=t_{\omega}^{*}$. Interestingly, the expression of $\psi_{\mathrm{incoh}}(\omega)$ in Eq. (31) is the same as for an illumination through the waveguide in Eq. (26), i.e. both incoherent two-photon amplitudes are equal if the external Bragg field drives the first atom with the same strengths as in the case where we probe the system through the waveguide. The coherent two-photon amplitude, however, differs strongly:
@@
Figure 7: Second-order correlation function at zero time-delay $g^{(2)}(0)$ under the Bragg condition as a function of $N$ for $\beta=0.01$ and two different detunings. In contrast to forward scattering, the point of perfect anti-bunching appears at $N=1$. For larger $N$ we recover $g^{(2)}(0)=1$. Interestingly, for $N=2$ an antibunching of $g^{(2)}(0)$ of 1/4 can be observed (black circle).
@@
$g^{(2)}(\tau)=\frac{|\left(\alpha_{1}^{\mathrm{sc}}\right)^{2}(1-t_{0}^{N})^{2}+g_{0}^{2}\psi_{\mathrm{incoh}}(\tau)|^{2}}{|\alpha_{1}^{\mathrm{sc}}(1-t_{0}^{N})|^{4}}.$ (33)
As expected, for $N=1$, one gets antibunching $g^{2}{(0)}=0$ (fluorescence from a single emitter) and for large $N$ the output light approaches $g^{2}{(0)}=1$ as shown in Fig. 7.
It is interesting to note that for $N=2$ we obtain $g^{(2)}(0)\simeq 1/4$. This is in contrast to the criteria of $g^{(2)}(0)\leq 1/2$ that is commonly used to verify the presence of a single quantum emitter. A $g^{(2)}(0)$ smaller than 1/2, in our case, is due to the constructive interference of the incoherent photons under Bragg illumination from different emitters. In contrast, if we consider $N$ emitters with shot-to-shot fluctuations of their relative distances, one recovers the common criteria $g^{(2)}(0)=1-1/N$ (see Appendix C). We note that $g^{(2)}(0)<1/2$ from two uncorrelated emitters has been already reported in the literature [74, 75].

### v2

@@ III.2 Single emitter, illumination through the waveguide
Now let us consider the case where the emitter is driven through the waveguide with Rabi frequency $\Omega_{\mathrm{in}}^{\mathrm{wg}}$, see Fig. 1, with the corresponding amplitude of the input coherent state $\alpha_{\mathrm{in}}=\Omega^{\mathrm{wg}}_{\mathrm{in}}/(2\sqrt{\beta\Gamma})$. The coherent component at the waveguide output $\alpha_{\mathrm{out}}$ now consists of the coherently scattered light $\alpha_{\mathrm{sc}}$ that interferes with the coherent field that is already present inside the waveguide $\alpha_{\mathrm{in}}$:
@@
In the limit of $\beta\ll 1$, the scattered field of a single atom is much smaller than the input field, $\alpha_{\mathrm{sc}}\ll\alpha_{\mathrm{in}}$, such that the two-photon wavefunction at the output is nearly identical to the one of the input, $\psi(\tau)\approx\alpha_{\mathrm{in}}^{2}$. Thus, a single weakly coupled emitter has a negligible effect on the photon statistics of the light transmitted in the waveguide. In contrast, as discussed in Sec. IV.1, the interaction of $N$ weakly coupled emitters, can strongly suppress or enhance the two-photon wavefunction at zero time delay $\psi(\tau=0)$.
@@ III.3 Single emitter, combined illumination
Let us now consider the case where the emitter is driven by both an external field and a guided field with Rabi frequencies $\Omega^{\mathrm{ext}}$ and $\Omega_{\mathrm{in}}^{\mathrm{wg}}$, respectively. Here and in the following, when considering combined illumination, for simplicity we assume the two fields in phase with identical polarizations and detunings, such that the Rabi frequencies sum up to the total Rabi frequency seen by the atom $\Omega=\Omega^{\mathrm{ext}}+\Omega_{\mathrm{in}}^{\mathrm{wg}}$. The combined scattering rate is then
@@
where we introduced the effective coupling constant $\beta^{\prime}=\beta\left(1+\Omega^{\mathrm{ext}}/\Omega_{\mathrm{in}}^{\mathrm{wg}}\right)$ which depends on the ratio of the driving field $\Omega_{\mathrm{in}}^{\mathrm{wg}}$ and the field resulting from the external Bragg beam $\Omega^{\mathrm{ext}}$. In analogy to Sec. III.2, one can introduce an effective transmission $t^{\prime}_{\Delta}$ and normalized scattering coefficient $g^{\prime}_{\Delta}$
@@
Note that in general $\beta^{\prime}$ is a complex number where $\mathrm{Arg}\{\beta\}$ is given by the relative phase between $\Omega^{\mathrm{ext}}$ and $\Omega^{\mathrm{wg}}$. In the following we restrict our discussions to the case where $\Omega^{\mathrm{ext}}$ and $\Omega^{\mathrm{wg}}$ are perfectly in phase such that $\beta^{\prime}$ is real and $\beta^{\prime}\geq 1$.
In summary, under a combined illumination, a single emitter effectively behaves, as if illuminated through the waveguide with an effectively increased coupling constant $\beta^{\prime}$ when the fields are in phase, as discussed in details in [69].
@@ IV Many emitters
Here, the coherent scattering amplitudes $\alpha_{n}^{\textrm{sc}}$ are computed according to Eq. (8) using the local Rabi frequencies $\Omega_{n}$ seen by the $n$-th atoms, which depend on the illumination configuration, see below. As previously, the coherent two-photon component is then given by $\psi_{\mathrm{coh}}=\alpha_{\mathrm{out}}^{2}$.
For deriving the incoherent two-photon component at the waveguide output, we resort to the frequency domain and find:
@@
where $\psi_{\mathrm{incoh}}^{(n)}(\omega)$ is the frequency-domain representation of the incoherent component of the two-photon amplitude of the $n$-th atom in the chain. We calculate it according to Eq. (11) using the local Rabi frequency $\Omega_{n}$, thereby assuming that the incoherent scattering process is only driven by the coherent component of the incident field. The product of amplitude transmission coefficients, $t_{\Delta+\omega}t_{\Delta-\omega}$, gives the amplitude transmission coefficient of a pair of incoherently scattered blue- and red-detuned photons past one atom. Note that, here, we neglect higher order nonlinear interaction where incoherently scattered photon pairs experience a second non-linear scattering. Neglecting these higher-order incoherent scattering processes is a good approximation in the regime of weak coupling ($\beta\ll 1$), see Sec. IV.1 where we explicitly compare our result with [34].
The output total two-photon amplitude in the time domain is thus given by $\psi(\tau)=\psi_{\mathrm{coh}}+\psi_{\mathrm{incoh}}(\tau)$. In the following, we apply this formalism to derive quantitative predictions for four experimentally interesting situations.
Figure 4: a) Contribution of the $n$-th atom to the incoherent two-photon component of the transmitted light. The first $n-1$ emitters attenuate the input light which is then incoherently scattered by the $n$-th emitter. Then, the following $N-n$ emitters act as a dispersive and absorptive medium for this incoherent two-photon component. b) The evolution of the coherent amplitude $\psi_{\mathrm{coh}}$ and incoherent photon amplitude $\psi_{\mathrm{incoh}}(\tau=0)$ at zero time delay as a function of atom number for two different detunings $\Delta$. The amplitudes are normalized to $\alpha_{\mathrm{in}}$. c) Second-order correlation function at zero time delay, $g^{(2)}(\tau=0)$, for different detunings. Plots are shown for $\beta=0.007$ as in [3].
@@ IV.1 N emitters, illumination through the waveguide
First, we discuss the case where $N$ atoms are solely excited through the waveguide mode. This setting was theoretically studied in [34] and [62] based on a scattering matrix formalism and higher order cumulant expansion, respectively, with a discussion of the resonant excitation case ($\Delta=0$). Here, we extend this discussion to off-resonant excitation, which has been experimentally studied and theoretically analyzed [3] using the model presented here.
@@
Inserting this expression into the incoherent two-photon component (see Eq. (11)) emitted by the $n$-th emitter into the waveguide, we obtain:
@@
Summing up the incoherent contribution of all emitters at the waveguide output (see Eq.  23), the incoherent contribution of all emitters, we obtain the incoherent component of the output two-photon wavefunction:
$\psi_{\mathrm{incoh}}(\omega)=(\Omega_{\mathrm{in}}^{\mathrm{wg}})^{2}\sum_{n=1}^{N}t_{\Delta}^{2(n-1)}\tilde{\psi}_{\mathrm{incoh}}(\omega)\left(t_{\Delta+\omega}t_{\Delta-\omega}\right)^{N-n}$
$=(\Omega_{\mathrm{in}}^{\mathrm{wg}})^{2}\,\tilde{\psi}_{\mathrm{incoh}}(\omega)\times\begin{cases}\frac{t_{\Delta}^{2N}-\left[t_{\Delta+\omega}t_{\Delta-\omega}\right]^{N}}{t_{\Delta}^{2}-t_{\Delta+\omega}t_{\Delta-\omega}}&\omega\neq 0\\ Nt_{\Delta}^{2(N-1)}&\omega=0\end{cases}$ (27)
@@
$\tilde{\psi}_{\mathrm{incoh}}(\omega)=\psi^{(n)}_{\mathrm{incoh}}(\omega)/\left(\Omega_{n}^{\mathrm{wg}}\right)^{2}~{}.$ (28)
As evident from Eq. 27, the incoherent component of the output two-photon wavefunction $\psi_{\mathrm{incoh}}(\omega)$ results from an interplay of (i) attenuation of coherent light along the ensemble, (ii) incoherent scattering at each atom and (iii) absorption of the incoherent photons by the following atoms, as sketched in Fig. 4.a). From the total two-photon wavefunction at the output, $\psi(\tau)$ , we obtain the second-order correlation function $g^{(2)}(\tau)$:
@@
Here, as $\psi_{\mathrm{incoh}}\propto\alpha_{\mathrm{in}}^{2}$, $g^{(2)}(\tau)$ is independent of the input power. This is a consequence of the truncation to the two-photon number state used here  [^7] Our theory is developed to leading order in $|\Omega|/\Gamma$ as in [34]. Correction to our model are on the order $P_{\mathrm{in}}\propto|\Omega|^{2}$ for $g^{2}(\tau)$, and $P_{\mathrm{}}^{2}$ for the squeezing spectrum.
@@
In this context, it is useful to define the phase difference between the incoherent and coherent component at zero time delay, $\Delta\varphi=\mathrm{Arg}\{\psi_{\mathrm{incoh}}(\tau=0)\}-\mathrm{Arg}\{\psi_{\mathrm{coh}}\}$. Depending on $\Delta\varphi$, the incoherent two-photon component will interfere with the coherent component either destructively (when $\Delta\varphi=\left(2m+1\right)\pi,\,m\in\mathbb{Z}$) or constructively (when $\Delta\varphi=2m\pi,\,m\in\mathbb{Z}$). This is illustrated in Fig. 4 b) where we show the evolution of both the coherent and incoherent two-photon amplitude in the complex plane as a function of the number of emitters for two different detunings. The coherent component still decays exponentially with the number of emitters, as in the case for resonant excitation. By comparison, the incoherent two-photon amplitude at zero time delay $\psi_{\mathrm{incoh}}(\tau=0)$ shows a more complex dynamics. Its magnitude first linearly increases due to a collective enhancement in forward scattering  [^8] See Appendix B and [2] for more details.. For larger atom number, the magnitude of the two-photon component decreases sub-exponentially with $N$ [34].
@@
Turning now to Fig. 4.c), it is apparent that off-resonant the combined dynamics of $\psi_{\mathrm{coh}}$ and $\psi_{\mathrm{incoh}}(\tau=0)$ leads to an oscillatory behavior of $g^{(2)}(\tau=0)$ as a function of $N$, which stems from the $N$-dependence of $\Delta\varphi$, see panel b), leading to alternating constructive and destructive interference of the two components. Surprisingly, when scanning $N$ and $\Delta$, we find that several combinations lead to perfect destructive interference and, hence, to photon antibunching (see orange and green lines in panel c). Specifically, this occurs when the moduli of both components are equal, $|\psi_{\mathrm{coh}}|=|\psi_{\mathrm{incoh}}(\tau=0)|$, and their phase difference is $\Delta\varphi=(2m+1)\pi$, $m\in\mathbb{Z}$. While the antibunching point on resonance ($\Delta=0$) has been predicted and observed in [34, 1], perfect antibunching for off-resonant excitations is a new prediction. It has been recently observed experimentally based on the predictions of this model [3].
We now compare our results to Mahmoodian et al. [34] where the two-photon wavefunction has been calculated using a scattering formalism together with the Bethe-Ansatz technique [65]. In contrast to our model, this approach yields exact solutions for any value of $\beta$. In Fig. 5 we compare the incoherent component $|\psi_{\mathrm{incoh}}(\tau=0)|$ calculated with the two approaches. Both predictions agree very well for $\beta\ll 1$. For larger $\beta$ deviations are expected as our model does not take into account higher order scattering processes. Note that, in contrast to [34], our model provides a closed expression for $\psi_{\mathrm{incoh}}(\tau=0)$, which is given by the frequency integral of the analytical expression given in Eq. (27). Thus, its numerical complexity does not depend on $N$, contrary to the numerical calculation based on [34]. Consequently, predictions made with our model are numerically much less expensive for large $N$.
@@ IV.2 External illumination: Bragg angle
where $\lambda_{f}=\lambda/n_{\mathrm{eff}}$ is the wavelength inside the waveguide, $n_{\mathrm{eff}}$ is the effective refractive index of the guided mode, and $a$ the lattice spacing and $m$ the Bragg order. This situation resembles the situation studied in [12, 74]. We will neglect free space interaction between the atoms such that we only consider Bragg conditions fulfilling $a\gtrsim\lambda$. We note that possible voids in the atomic array would not affect our predictions.
For a plane wave, each emitter is subject to an identical external driving field with a Rabi frequency of magnitude $|\Omega^{\mathrm{ext}}|$. The first emitter scatters a coherent amplitude $\alpha_{1}^{\mathrm{sc}}$ into the waveguide which constructively interferes with the scattered amplitude from the second emitter according to the Bragg condition of Eq. (30). After $N$ emitters, and taking into account absorption, the output two-photon coherent amplitude inside the waveguide is
$\alpha_{\mathrm{out}}$ $=\sum_{n=1}^{N}t_{0}^{N-n}\alpha_{1}^{\mathrm{sc}}=\frac{1-t_{0}^{N}}{1-t_{0}}\alpha_{1}^{\mathrm{sc}}=\frac{1-t_{0}^{N}}{2\beta}\alpha_{1}^{\mathrm{sc}}.$ (31)
@@
using $t_{-\omega}=t_{\omega}^{*}$. Interestingly, the expression of $\psi_{\mathrm{incoh}}(\omega)$ in Eq. (32) is the same as for an illumination through the waveguide in Eq. (27) for $\Delta=0$, i.e. both incoherent two-photon amplitudes are equal if the external Bragg field drives the first atom with the same strength as in the case where we probe the system through the waveguide. The coherent two-photon amplitude, however, differs strongly, and for external illumination we obtain:
@@
Figure 7: Second-order correlation function at zero time-delay $g^{(2)}(0)$ under the Bragg condition as a function of $N$ for $\beta=0.01$. As in the excitation scheme with a single emitter from the side (Sec. III.1), the point of perfect anti-bunching appears at $N=1$. For larger $N$, $g^{(2)}(0)$ approaches 1. Interestingly, for $N=2$ an antibunching of $g^{(2)}(0)$ of 1/4 is predicted (black circle).
@@
$g^{(2)}(\tau)=\frac{|\left(\alpha_{1}^{\mathrm{sc}}\right)^{2}(1-t_{0}^{N})^{2}+4\beta^{2}\psi_{\mathrm{incoh}}(\tau)|^{2}}{|\alpha_{1}^{\mathrm{sc}}(1-t_{0}^{N})|^{4}}.$ (34)
As expected, for $N=1$, one gets antibunching $g^{(2)}{(0)}=0$ (fluorescence from a single emitter) and for large $N$ the output light approaches $g^{(2)}{(0)}=1$ as shown in Fig. 7.
It is interesting to note that for $N=2$ we obtain $g^{(2)}(0)\simeq 1/4$. This is in contrast to the criteria of $g^{(2)}(0)\leq 1/2$ that is commonly used to verify the presence of a single quantum emitter. A $g^{(2)}(0)$ smaller than 1/2, in our case, is due to the constructive interference of the incoherent photons under Bragg illumination from different emitters. In contrast, if we consider $N$ emitters with shot-to-shot fluctuations of their relative distances, one recovers the common criteria $g^{(2)}(0)=1-1/N$ (see Appendix C). We note that $g^{(2)}(0)<1/2$ from two coherently emitters has been predicted in the literature [76, 77].